# Learning finitely correlated states: stability of the spectral reconstruction

### v2

@@ 5 Non-homogeneous FCS on a finite chain
Definition 5.1 (Finitely correlated states on a finite chain).
@@
and $\Omega^{[i,j,k]}=\Omega^{[\max\{1,i\},j,\min\{k,N\}]}$ for all the other values of $i$ and $k$, with $0\leq j\leq N$. A state $\omega\in\mathcal{A}_{[1,N]}$ is $(r,l,m)$ finitely correlated if $\mathrm{rank}\,\Omega^{[1,j,N]}=\mathrm{rank}\,\Omega^{[j-l+1,j,j+r]}=:m_{j}\leq m$ for $0\leq j\leq N$. We also use the notation $\Omega^{[i,j,r]}_{X}:=\Omega^{[i,j,r]}(X\otimes\cdot)$.
@@
Proposition 5.2 (Observable realization).
@@
The proof is given in the appendix C.1. Note that $m_{N}=m_{0}=1$ and thus we do not need the analogue of $e$ and $\rho$ in this case because they are essentially included in $\mathbb{K}^{(N)}:\mathbb{C}\otimes\mathcal{A}\to\mathbb{C}^{m_{N-1}}$ and $\mathbb{K}^{(1)}:\mathbb{C}^{m_{1}}\otimes\mathcal{A}\to\mathbb{C}$ respectively. In the following, we use the notation $\mathbb{K}^{[i,j]}:=\mathbb{K}^{(i)}\mathbb{K}^{(i+1)}\cdots\mathbb{K}^{(j)}$.
Proposition 5.3 (Empirical realizations).
@@
A similar reconstruction algorithm can be run in this case, where $\hat{\Omega}^{[j-l,j,j+r]}$ is an estimate of ${\Omega}^{[j-l,j,j+r]}$, $\hat{\Omega}_{(\cdot)}^{[j-l-1,j-1,j+r]}$ is an estimate of ${\Omega}_{(\cdot)}^{[j-l-1,j-1,j+r]}$ and $\hat{U}_{j}$ is obtained from the $m_{j}$-truncated singular value decomposition of $\hat{\Omega}^{[j-l,j,j+r]}$.
Definition 5.4 (Spectral state reconstruction, non-homogeneous case).
@@
and $\hat{U}_{j}\hat{D}_{j}\hat{W}_{j}^{\intercal}$is the $m_{j}$-truncated singular value decomposition of $\hat{\Omega}^{[l-j+1,j,l+r]}$.
@@
We then have
Theorem 5.5 (Error propagation for general orders and finite-size chain).
For the state reconstruction in Definition 4.1, and the error parameter defined in Eq. (157), if $\hat{U}_{j}^{\intercal}U_{j}$ are invertible, then
@@ 5.1 Bound for $(r,l,m)$-FCS
$\|B\|_{e_{j+1}\rightarrow e_{j},cb}:=\sup\{\|(\mathrm{id}_{\mathbb{M}_{n}}\otimes B)(X)\|_{e_{j}^{(n)}}\ |\ X\in\mathbb{C}^{m_{j}}\otimes\mathbb{M}_{n},\|X\|_{e_{j+1}^{(n)}}\leq 1\}.$ (170)
On the other hand, as we did in Section 3, we can also define norms from the inner product defined by the maps $\Omega^{[j-l,j,j+r]}$, observing that any vector $v$ in $\mathbb{C}^{m_{j}}$ can be written as $\mathbb{K}^{[j+1,N]}_{X(v)\otimes{\mathds{1}}^{\otimes N-j-r}}$ for some $X(v)\in\mathcal{A}_{[j+1,j+r]}$, and then $(v,w)_{\Omega^{[j-l,j,j+r]}}:=\sum_{i=1}^{d_{\mathcal{A}}^{l}}\omega(Y_{i}\otimes X(v))\omega(Y_{i}\otimes X(w)),$ with $\{Y_{i}\}$ being a self-adjoint orthonormal basis of $\mathcal{A}_{[j-l,j]}$. We denote such norms as $\|\cdot\|_{2,j}$, their amplifications as $\|\cdot\|_{2,j,n}$ and by the same argument of Lemma 3.1 and Remark 3.2 and therefore, as in (51), we have the inequalities
@@
By using the error progapagation bound in Theorem 5.5, we obtain
Theorem 5.6.
Let $\omega$ be a $(l,r,m)$-finitely correlated state, with associated map $\Omega$. There exists an algorithm which takes as input $n$ copies of $\omega$ and a threshold $\eta$, such that it makes measurements on marginals of size $l+r+1$ and it outputs an estimate $\hat{\omega}$ such that, with high probability,
$d_{\mathrm{Tr}}(\hat{\omega},\omega)\leq\epsilon,$ (173)
whenever $\min_{j=0,\dots,N}\sigma_{m}(\Omega^{[j-l+1,j,j+r]})/2\geq\eta\geq 2\varepsilon$ with $\varepsilon=\min_{j=1,...,N}\frac{\epsilon\sigma^{3}_{m}(\Omega^{[j-l+1,j,j+r]})}{20tm\sqrt{d_{\mathcal{A}}}}$ and $n$ larger than the number of copies of $\omega$ necessary for learning the marginals $\omega_{[\max(j-l+1,1),\min(j+r,N)]}$ and $\omega_{[\max(j-l,1),\min(j+r,N)]}$, $j=0,...,N$ in Hilbert-Schmidt distance at precision $\varepsilon$.
5.2 Bound for $C^{*}$-$(r,l,m)$-FCS
We can also define realizations for states on a finite chain; note that any state on a finite chain admits a realization of this kind, but the dimension of the memory system vector space can grow exponentially with $N$ in general.
Definition 5.7.
A $C^{*}$-realization of a state $\omega$ is defined by a unital $C^{*}$-algebra $\mathcal{B}$, a state $\rho_{0}\in\mathcal{B}^{*}$ and a collection of maps $\mathcal{E}^{(j)}:\mathcal{A}\otimes\mathcal{B}\rightarrow\mathcal{B}$ such that
@@ 5.2 Bound for $C^{*}$-$(r,l,m)$-FCS
This representation suggests that the state is generated by sequential maps on a quantum memory system. Note that in favour of a simpler presentation we are choosing a (suboptimal) memory system that does not depend on $j$. If we allowed the memory system to change with $j$ we could absorb $\rho_{0}$ and $\mathds{1}_{B}$ in the definitions of respectively $\mathcal{E}^{(1)}$ and $\mathcal{E}^{(N)}$ as in Proposition 5.2.
@@
$\mathcal{W}_{j}$ =$=\{\mathcal{E}^{[j+1,j+r]}(A\otimes\mathds{1}_{\mathcal{B}}),A\in\mathcal{A}_{[j+1,j+k]}\},$ (177)
$\tilde{\mathcal{W}}_{j}$ =$=\{\rho_{j-l+1}\mathcal{E}^{[j-l+1,j]}(A\otimes\cdot),A\in\mathcal{A}_{[j-l+1,j]}\}.$ (178)

### v3

@@ Appendix D Non-translation-invariant FCS on a finite chain
Definition D.1 (Finitely correlated states on a finite chain).
@@
and $\Omega^{[i,j,k]}=\Omega^{[\max\{1,i\},j,\min\{k,N\}]}$ for all the other values of $i$ and $k$, with $0\leq j\leq N$. A state $\omega\in\mathcal{A}^{*}_{[1,N]}$ is $(r,l,m)$ finitely correlated if $\mathrm{rank}\,\Omega^{[1,j,N]}=\mathrm{rank}\,\Omega^{[j-l+1,j,j+r]}=:m_{j}\leq m$ for $0\leq j\leq N$. We also use the notation $\Omega^{[i,j,k]}_{X}:=\Omega^{[i,j,k]}(X\otimes\cdot)$.
@@
Proposition D.2 (Observable realization).
@@
Proof.
Let $V_{[i+1,j]}$ be the space generated by the span of the functionals $\Psi^{(i+1,j)}_{X}\in\mathcal{A}_{[\max(1,i-l+1),\min(i,N)]}^{*}$, $X\in\mathcal{A}_{[\max(1,i+1),\min(j,N)]}$, $\Psi^{(i+1,j)}_{X}(Y):=\omega(Y\otimes X)$ (where $V_{[N+1,N]}$ is one-dimensional and it is the span of the functional $\Psi(Y)=\omega(\mathds{1}\otimes Y)$, $Y\in\mathcal{A}_{[N-l,N]}$, and $V_{[1,j]}$ is one-dimensional as it is the span of functionals $\Phi(X)=\omega(X\otimes\mathds{1})$, $X\in\mathcal{A}_{[1,j]}$ also a one-dimensional vector space). For $i=1,...,N$ we can define the linear maps $\mathbb{E}_{A}^{(i)}:V_{[i+1,i+r+1]}\rightarrow V_{[i,i+r]}$ as $\mathbb{E}^{(i)}_{A}\Psi^{(i+1,i+r+1)}_{X}=\Psi^{(i,i+r+1)}_{A\otimes X}=\Psi^{(i,i+r)}_{X^{\prime}}$, where the last equality holds for some $X^{\prime}\in A_{[\max(1,i+1),\min(N,i+r)]}$ because of the definition of $(l,r,m)$-finitely correlated state. We obtain
The image of $\Omega^{[j-l+1,j,j+r]}$ lies in $V_{[j+1,j+r+1]}$, so we can write
where $\Omega^{[j-l,j-1,j+r]}_{X_{j}}:\mathcal{A}_{[j,\min(j+r,N)]}\rightarrow V_{[j,j+r]}=V_{[j,j+r-1]}$, $\Omega^{[j-l,j-1,j+r]}_{X_{j}}(X)=\Psi^{(j,j+r)}_{X_{j}\otimes X}$. Using the singular value decomposition
$\Omega^{[j-l+1,j,j+r]}={{U}}_{j}D_{j}O_{j},$ (183)
with $U_{j}:\mathbb{R}^{m_{j}}\rightarrow V_{[j+1,j+r+1]}$, $D:\mathbb{R}^{m_{j}}\rightarrow\mathbb{R}^{m_{j}}$, $O_{j}:\mathcal{A}_{[j+1,j+r]}\rightarrow\mathbb{R}^{m_{j}}$. We get
$\mathbb{E}_{X_{N}}^{(N)}$ $=\Omega^{[N-l,N-1,N+r]}_{X_{N}}$ (184)
Note that for $X\in A_{[j+1,j+r]}$, ${U}_{j}{U}_{j}^{\intercal}\Psi^{[j+1,j+r]}_{X}={U}_{j}{U}_{j}^{\intercal}\Omega^{[j-l+1,j,j+r]}(X)=\Omega^{[j-l+1,j,j+r]}(X)=\Psi^{[j+1,j+r]}_{X}$, therefore we have
Note that $m_{N}=m_{0}=1$ and thus we do not need the analogue of $e$ and $\rho$ in this case because they are essentially included in $\mathbb{K}^{(N)}:\mathbb{C}\otimes\mathcal{A}\to\mathbb{C}^{m_{N-1}}$ and $\mathbb{K}^{(1)}:\mathbb{C}^{m_{1}}\otimes\mathcal{A}\to\mathbb{C}$ respectively. In the following, we use the notation $\mathbb{K}^{[i,j]}:=\mathbb{K}^{(i)}\mathbb{K}^{(i+1)}\cdots\mathbb{K}^{(j)}$.
Proposition D.3 (Empirical realizations).
@@
A reconstruction algorithm similar to Algorithm 1 can be run in this case fixing a threshold $\eta$, with inputs $\hat{\Omega}^{[j-l,j,j+r]}$ as an estimate of ${\Omega}^{[j-l,j,j+r]}$, $\hat{\Omega}_{(\cdot)}^{[j-l-1,j-1,j+r]}$ as an estimate of ${\Omega}_{(\cdot)}^{[j-l-1,j-1,j+r]}$ and computing $\hat{U}_{j}$ from obtained from the singular value decomposition of $\hat{\Omega}^{[j-l,j,j+r]}$ truncated to the singular values larger than $\eta/2$. As in the translation invariant case, as soon as the Hilbert-Schmidt norm errors are less than $\eta/3$, the matrices $\hat{U}_{j}$ correspond to those obtained from truncation to the first $m_{j}$ singular values. We refer to this algorithm as LearnFCS(l,r). The estimated realization that it aims to computes is the following:
Definition D.4 (Spectral state reconstruction, non-homogeneous case).
@@
and $\hat{U}_{j}\hat{D}_{j}\hat{O}_{j}$ is the singular value decomposition of $\hat{\Omega}^{[j-l+1,j,j+r]}$ truncated to the first $m_{j}$ singular values.
@@
We also denote $\hat{M}_{j}\coloneqq\hat{U}^{\intercal}_{j}U_{j}M_{j}$. We then have
Theorem D.5 (Error propagation for general orders and finite-size chain).
For the state reconstruction in Definition C.1, and the error parameter defined in Eq. (195), if $\hat{U}_{j}^{\intercal}U_{j}$ are invertible, then
@@ D.1 Bound for $(r,l,m)$-FCS
$\|B\|_{e_{j+1}\rightarrow e_{j},cb}:=\sup\{\|(\mathrm{id}_{\mathbb{M}_{n}}\otimes B)(X)\|_{e_{j}^{(n)}}\ |\ X\in\mathbb{C}^{m_{j}}\otimes\mathbb{M}_{n},\|X\|_{e_{j+1}^{(n)}}\leq 1,n\in\mathbb{N}\}.$ (208)
On the other hand, as we did in Section B, we can also define norms from the inner product defined by the maps $\Omega^{[j-l,j,j+r]}$, observing that any vector $v$ in $\mathbb{C}^{m_{j}}$ can be written as $\mathbb{K}^{[j+1,N]}_{X(v)\otimes{\mathds{1}}^{\otimes N-j-r}}$ for some $X(v)\in\mathcal{A}_{[j+1,j+r]}$, and then $(v,w)_{\Omega^{[j-l,j,j+r]}}:=\sum_{i=1}^{d_{\mathcal{A}}^{2l}}\omega(Y_{i}\otimes X(v))\omega(Y_{i}\otimes X(w)),$ with $\{Y_{i}\}$ being a self-adjoint orthonormal basis of $\mathcal{A}_{[j-l,j]}$. We denote such norms as $\|\cdot\|_{2,j}$, their amplifications as $\|\cdot\|_{2,j,n}$ and by the same argument of Lemma B.6 and Remark B.7 and therefore, as in (82), we have the inequalities
@@
By using the error progapagation bound in Theorem D.5, we obtain
Theorem D.6.
Let $\omega$ be a $(l,r,m)$-finitely correlated state, with associated maps $\Omega^{[i,j,k]}$ satisfying $\eta\leq\min_{j=0,\dots,N}\sigma_{m}(\Omega^{[j-l+1,j,j+r]})/2$. Assume we have estimates of marginals $\omega_{[\max(j-l,1),\min(j+r,N)]}$, $\omega_{[\max(j-l+1,1),\min(j+r,N)]}$, $j=0,...,N$, such that the Hilbert-Schmidt norm errors are smaller than $\frac{\epsilon\eta^{3}}{20tm\sqrt{d_{\mathcal{A}}}}$. The estimated realization parameters from LearnFCS(l,r) (given by D.4) give rise to an estimated operator $\hat{\omega}$ such that
$\frac{\|\hat{\omega}-\omega\|_{1}}{2}\leq\epsilon,$ (211)
A sample complexity bound follows immediately as in the non-translation invariant case.
D.2 Bound for $C^{*}$-$(r,l,m)$-FCS
We can also define $C^{*}$-realizations for states on a finite chain; note that any state on a finite chain admits a realization of this kind, but the dimension of the memory system vector space can grow exponentially with $N$ in general.
Definition D.7.
A $C^{*}$-realization of a state $\omega$ is defined by a unital $C^{*}$-algebra $\mathcal{B}$, a state $\rho_{0}\in\mathcal{B}^{*}$ and a collection of unital maps $\mathcal{E}^{(j)}:\mathcal{A}\otimes\mathcal{B}\rightarrow\mathcal{B}$ such that
@@ D.2 Bound for $C^{*}$-$(r,l,m)$-FCS
This representation suggests that the state is generated by sequential maps on a quantum memory system. Note that in favour of a simpler presentation we are choosing a (suboptimal) memory system that does not depend on $j$. If we allowed the memory system to change with $j$ we could absorb $\rho_{0}$ and $\mathds{1}_{B}$ in the definitions of respectively $\mathcal{E}^{(1)}$ and $\mathcal{E}^{(N)}$ as in Proposition D.2.
@@
$\mathcal{W}_{j}$ =$=\{\mathcal{E}^{[j+1,\min(j+r,N)]}(A\otimes\mathds{1}_{\mathcal{B}}),A\in\mathcal{A}_{[j+1,\min(j+r,N)]}\},$ (215)
$\tilde{\mathcal{W}}_{j}$ =$=\{\rho_{\max(j-l+1,1)}\mathcal{E}^{[\max(j-l+1,1),j]}(A\otimes\cdot),A\in\mathcal{A}_{[\max(j-l+1,1),j]}\}.$ (216)